OpenStudy (anonymous):

A pound of nails cost $2. A pound of screws cost $6. How many pounds of nails are needed for a 5–pound mixture that costs $3.60 per pound?

OpenStudy (anonymous):

Let n and s equal the weight in lbs of the nails and screws respectively. Solve the following two equations for n and s. The first equation relates the costs in dollars and the second equation relates the weights in pounds.\[\{2n + 6s=5 *3.6,n+s=5\}\]\[{n = 3, s = 2} \]